OpenStudy (yrivers36):

Directions says use technology to solve 0.1x+0.2y=50 0.3x-0.1y=10

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=solve+0.1x%2B0.2y%3D50+%2C0.3x-0.1y%3D10 wolfram|alpha counts as technology

jhonyy9 (jhonyy9):

multiplí both sides with10 on every two equation and will get x+2y=500 3x+y=100

jhonyy9 (jhonyy9):

x+2y=500 3x+y=100 /*2 multiply with 2 and subtract from first equation

jhonyy9 (jhonyy9):

x+2y=500 6x+2y=200 ----------- -5x 0 =300 -5x=300 x=300/-5 = - 60 -60+2y=500 2y=560 y=560/2=280 y=280 x= -60 y= 280